# Deductive Beam Search: Decoding Deducible Rationale for Chain-of-Thought Reasoning

### v2

@@ Abstract
Recent advancements have significantly augmented the reasoning capabilities of Large Language Models (LLMs) through various methodologies, especially chain-of-thought (CoT) reasoning. However, previous methods fail to address reasoning errors in intermediate steps, leading to accumulative errors. In this paper, we propose Deductive Beam Search (DBS), which seamlessly integrates CoT and deductive reasoning with step-wise beam search for LLMs. Our approach deploys a verifier, verifying the deducibility of a reasoning step and its premises, thus alleviating the error accumulation. Furthermore, we introduce a scalable and labor-free data construction method to amplify our model’s verification capabilities. Extensive experiments demonstrate that our approach significantly enhances the base performance of LLMs of various scales (7B, 13B, 70B, and ChatGPT) across 8 reasoning datasets from 3 diverse reasoning genres, including arithmetic, commonsense, and symbolic. Moreover, our analysis proves DBS’s capability of detecting diverse and subtle reasoning errors and robustness on different model scales [^1] Code is available at https://github.com/OSU-NLP-Group/Deductive-Beam-Search..
Machine Learning, ICML
@@ 1 Introduction
Machine reasoning has witnessed tremendous progress thanks to the emergence of Large Language Models (LLMs) (OpenAI, 2023; Google, 2023a, b; Meta, 2023). The power of LLMs activates the ability to conduct step-by-step chain-of-thought (CoT) reasoning (Wei et al., 2022b, a), significantly boosting the performance of reasoning tasks (Wang et al., 2022; Paul et al., 2023; Lyu et al., 2023).
Although CoT reasoning has demonstrated the superiority of step-by-step reasoning, its dependency on intermediate steps inevitably introduce accumulative errors (Du et al., 2023; Yu et al., 2023) in the process, as shown in Figure 1. Previous research that alleviates these errors lies in two main paradigms: 1) Answer aggregation across multiple rationales. They utilize majority voting (Wang et al., 2022) or deploy a verifier to score on each rationale (Li et al., 2023). However, these methods do not directly address errors in the reasoning process, undermining the reliability of their outcomes. 2) Intermediate step correction. This line of works decomposes the reasoning path into reasoning steps and applies self-correction on each step (Weng et al., 2023; Ling et al., 2023; Paul et al., 2023; Xie et al., 2023). Yet, recent research finds that, without external feedback, LLMs tend to modify reasoning steps regardless of their correctness (Huang et al., 2023; Hong et al., 2023).
@@
Previous works fail to address reasoning errors in intermediate steps, compromising the ability to conduct systematic reasoning. To mitigate this issue, we embrace the principle of deductive reasoning (Clark, 1969; Johnson-Laird, 1999, 2010) to navigate CoT. In deductive reasoning, every step logically follows its premises, where a deducible reasoning step is termed a logical consequence (Dinkmeyer, 1976; Hanson, 1997). A key attribute of logical consequence is that if the premises hold, the deducible reasoning step is true, suggesting a correct outcome. Inspired by this attribute, we propose to navigate CoT towards a more deducible path.
@@
Nonetheless, challenges arise when introducing the principle of deductive reasoning into CoT reasoning without changing the standard prompt paradigm and the parameters of LLMs. 1) Navigation on CoT reasoning. Since LLMs cannot always conduct correct deductive reasoning, they have to explore the potential reasoning space and then choose reasoning steps that are more likely to be deducible, which brings the trade-off between exploration and exploitation (Donoso et al., 2014; Dasgupta et al., 2019). 2) Verification of deducibility. On one hand, previous research shows it is hard for LLMs to detect reasoning errors (Hong et al., 2023). On the other hand, symbolic reasoning engines (Cavada et al., 2014; Li et al., 2018) can reliably verify the correctness. However, transferring natural language to symbolic language without losing generality remains an unsolved problem in machine reasoning.
@@
Confronted with these challenges, we propose Deductive Beam Search (DBS), adaptable to all models and settings, as shown Figure 2. For the trade-off challenge, we decompose the reasoning process into reasoning steps and incorporate step-wise beam search. In terms of the verification challenge, we propose a deductive verifier, which takes a reasoning step and its premises as inputs and outputs a deductive score, evaluating the logical coherence between them. Specifically, LLM samples a list of potential reasoning steps to explore. Then, our deductive verifier exploits by selecting steps that are more deducible. To train an effective verifier, we propose a scalable way of synthesizing fine-grained and diverse deductive reasoning errors without human annotation. Initially, the verifier is trained to verify heuristically synthesized wrong steps with typical reasoning error patterns. Subsequently, we ask LLMs to generate reasoning steps where false ones detected by our verifier serve as hard negatives. These hard negatives are adopted to train a deductive verifier with model feedback.
@@ 2.1 Multi-Step Chain-of-Thought Reasoning
$\Pr{}_{LM}(\textbf{r}_{[t]}|\textbf{q})$ $=\Pr{}_{LM}(\textbf{r}_{1}|\textbf{q})\times\prod_{i=1}^{t-1}{\Pr{}_{LM}(\textbf{r}_{i+1}|\textbf{q},\textbf{r}_{[i]})}$ (2)
$=\prod_{i=0}^{t-1}{\Pr{}_{LM}(\textbf{r}_{i+1}|\textbf{q},\textbf{r}_{[i]})}.$
@@ 2.2 Step-wise Beam Search
$Y_{[i]}=\operatorname*{arg\,max}_{\textbf{y}_{[i]}^{1},\textbf{y}_{[i]}^{2},...,\textbf{y}_{[i]}^{m}}\sum_{k\in[m]}\Phi(\textbf{y}_{[i]}^{k}).\vspace{0.0005cm}$ (3)
@@
$R_{[i]}=\operatorname*{arg\,max}_{\textbf{r}_{[i]}^{1},\textbf{r}_{[i]}^{2},...,\textbf{r}_{[i]}^{m}}\sum_{k\in[m]}\Psi(\textbf{r}_{[i]}^{k}).\vspace{0.0005cm}$ (4)
Combining multi-step CoT reasoning with step-wise beam search balances exploration and exploitation in reasoning tasks. However, confidence scores from language models cannot verify logical consequence between a reasoning step and its premises.
Table 1: Examples of heuristically synthesized false reasoning steps. The notations correspond to the notations in Section 3.
Context Type Not. Reasoning Step Error Reason
Randy has some money. He spent \$10 buying his lunch. He spent a quarter of the money he had left on an ice cream cone. If the ice cream cone cost \$5, what is the amount of money, in dollars, Randy had at first? Gold r
Randy has 5*4=20 dollars left after buying lunch.
Grounding $\textbf{r}_{1}^{\prime}$
Randy has 10*4=40 dollars left after buying lunch.
Minor token-level error, hard for models to detect.
Logic $\textbf{r}_{2}^{\prime}$
At first, Randy had a sum of 20+10=30 dollars.
Logic-level error caused by reversed steps, not following deductive reasoning.
Irrelevant $\textbf{r}_{3}^{\prime}$ He eats 65 black cookies from the cookie jar, with 1/2 * 130 = 65.
Major error, completely incoherent with the context.
To tackle this problem, we propose to constrain the step-wise beam search with deductive scores.
@@ 2.3 Deductive Verification Constrained Beam Search
$s_{j}=\textbf{s}([\textbf{r}_{[i-1]},\textbf{r}_{j}])$ $=\textbf{s}(\textbf{r}_{[i-1]})\times\Pr{}_{f}(\textbf{r}_{j}|\textbf{q},\textbf{r}_{[i-1]})$ (5)
$=\prod_{k=1}^{i}{\Pr{}_{f}(\textbf{r}_{k}|\textbf{q},\textbf{r}_{[k-1]}}),$
@@ 3 Deductive Verifier
As stated above, a deductive verifier evaluates whether the reasoning step can be deduced from previous contexts, which resembles a natural language inference (NLI) task. Thus, we use deberta-v3-large (He et al., 2023), which achieves the best performance across various NLI benchmarks despite its small size, as the backbone. A small scalar head is adopted to predict deductive scores based on embedding the [CLS] token.
However, the difficulties of training a deductive verifier primarily reside in the training data quality and the training method. Changing one single token could lead to various errors, which is hard for any model to detect. Furthermore, the lack of high-quality false deductive reasoning step hinges on the training of the verifier. To fully understand how LLMs make mistakes, we dive into the incorrect samples generated by LLMs. From the perspective of deductive reasoning, there are two main classes of reasoning errors: grounding errors and logic errors (Ling et al., 2023). Most grounding errors happening in the reasoning process can be detected by finding the contradiction between the context given and the generated rationales, while the latter ones are illogical reasoning steps deduced from the previous context.
@@ 3.1 A General Deductive Verifier
where $m_{i}$ is the hyper-parameter controlling the margin and $\alpha_{i}$ is the hyper-parameter weighing each loss.
@@ 3.2 Deductive Verifier with Model Feedback
In the first stage, we train a general deductive verifier, but the wrong samples synthesized heuristically are less diverse than the ones encountered during inference. To bridge the gap between synthesized data and real-world data, we use the verifier from stage 1 to detect false reasoning steps generated by an actual language model, where we choose Llama2-7b for the generation. The reason why we choose a relatively small language model for the generation is to maximize the diversity and the likelihood of generating incorrect reasoning steps.
To be concrete, given the verifier $f_{1}$ trained by stage 1, we feed current context c into the language model and sample 10 reasoning steps. Then, these reasoning steps are scored and ranked by $f_{1}$. From this ranking, we select the reasoning step that exhibits the most significant decrease in the deductive score, designating it as the hard negative sample. We replace $\textbf{r}_{1}^{\prime}$ with the generated hard negative sample as $\textbf{r}_{1}^{\prime\prime}$, keeping the original way of generating $\textbf{r}_{2}^{\prime}$ and $\textbf{r}_{3}^{\prime}$. Consequently, we continue training the verifier $f_{1}$ by Equation 6 with a smaller learning rate.
Table 2: The result comparison on arithmetic reasoning, commonsense reasoning, and symbolic reasoning tasks. The results represent accuracy (%) on each dataset. Bold indicates best results and underline indicates second bests. The white area represents single reasoning chain setting, where greedy represents the greedy decoding algorithm (Jurafsky & Martin, ). The grey area represents multiple reasoning chains, where we decode 10 reasoning paths for each method. SC represents self-consistency (Wang et al., 2022).
@@
4 Experimental Setup
4.1 Reasoning Tasks
Symbolic Reasoning. We use the Coin Flip dataset from BIG-bench (bench authors, 2023). The task is to determine which face of the coin is up after a series of operations.
4.2 Details
Language Models. We evaluate our method on models of various scales, including Llama2-7b, Llama2-13b, Llama2-70b (Meta, 2023), and ChatGPT (gpt-3.5-turbo-instruct) (OpenAI, 2022). These models represent different levels of reasoning abilities.
Verifier. We choose deberta-v3-large as the backbone of our verifier. At stage 1, the general verifier is trained on 22,363 samples generated from the GSM8K dataset. At stage 2, the verifier for arithmetic and symbolic reasoning is trained on 150,000 samples generated from MetaMathQA, while the verifier for commonsense reasoning is trained on 5,000 samples generated from StrategyQA. The training details are in Appendix B.1 and B.2.
@@ 5 Main Result
Table 2 demonstrates the overall performance of the methods. We compare DBS with baselines under two paradigms: single reasoning chain setting and multiple reasoning chain setting. In multiple reasoning chain setting, the generated outcomes are integrated with self-consistency.
Table 3 presents a comparative analysis of our approach against other decoding strategies through a self-evaluate pattern (Xie et al., 2023). Since the self-evaluate pattern is token-expensive, we only evaluate the methods on 100 randomly sampled questions from the GSM8K dataset on Llama2-7b and GPT-3.5, which are two typical examples of models with distinct levels of abilities.
Table 3: Comparison of different decoding methods and evaluation methods. We test all the listed methods on 100 randomly sampled arithmetic reasoning questions from the GSM8K dataset. We report accuracy, output tokens (including evaluation tokens used by other methods), and the number of rationales for self-consistency.
@@
Llama2-7b SelfEval 24 5,371,280 1
SelfEval + SC 22 10
DBS 38 297,301 1
DBS + SC 43 10
ChatGPT SelfEval 70 1,854,147 1
SelfEval + SC 74 10
DBS 81 474,221 1
DBS + SC 85 10
5.1 Effectiveness
As shown in Table 2, DBS improves the performance across models of different scales and diverse reasoning tasks. For the single reasoning chain setting, the improvement is substantial. On arithmetic reasoning tasks, taking GSM8K as an example, we observe an increase from 7.6% to 16.6% across models of various scales. Specifically, with Llama2-7b and gpt-3.5-turbo-instruct, DBS yields improvements of 9.2% and 7.0%, respectively, affirming the effectiveness of our proposed decoding strategy. On commonsense reasoning tasks and symbolic reasoning tasks, we can expect an average increase of 2.5%/2.0% on models of all scales.
5.2 Comparison with Decoding Strategies
Table 3 compares our decoding strategy with previous SOTA decoding strategies. The comparative results, grounded in accuracy and token cost metrics, substantiate our approach’s effectiveness and token efficiency. Notably, the self-evaluate pattern consumes an excessive amount of tokens during evaluation. In contrast, DBS enhances the performance by approximately 10% on both Llama2-7b and gpt-3.5-turbo-instruct, with fewer tokens generated across the paradigms of single reasoning chain and multiple reasoning chains.

### v3

@@ Abstract
Recent advancements have significantly augmented the reasoning capabilities of Large Language Models (LLMs) through various methodologies, especially chain-of-thought (CoT) reasoning. However, previous methods often struggle to address reasoning errors in intermediate steps, which can lead to accumulative errors. In this paper, we propose Deductive Beam Search (DBS), which seamlessly integrates CoT and deductive reasoning with step-wise beam search for LLMs. Our approach deploys a verifier, verifying the deducibility of a reasoning step and its premises, thus alleviating the error accumulation. Furthermore, we introduce a scalable and labor-free data construction method to amplify our model’s verification capabilities. Extensive experiments demonstrate that our approach significantly enhances the base performance of LLMs of various scales (7B, 13B, 70B, and ChatGPT) across 8 reasoning datasets from 3 diverse reasoning genres, including arithmetic, commonsense, and symbolic. Moreover, our analysis proves DBS’s capability of detecting diverse and subtle reasoning errors and robustness on different model scales. Data and codes are released at https://github.com/OSU-NLP-Group/Deductive-Beam-Search.
@@ 1 Introduction
Machine reasoning has witnessed tremendous progress thanks to the emergence of Large Language Models (LLMs) (OpenAI, 2023; Google, 2023; Anil et al., 2023; Touvron et al., 2023; McIntosh et al., 2023). The power of LLMs activates the ability to conduct step-by-step chain-of-thought (CoT) reasoning (Wei et al., 2022b; a), significantly boosting the performance of reasoning tasks (Wang et al., 2022; Paul et al., 2023; Lyu et al., 2023).
@@
Although CoT reasoning has demonstrated the superiority of step-by-step reasoning, its dependency on intermediate steps inevitably introduces accumulative errors (Du et al., 2023; Yu et al., 2023; Wu et al., 2024) in the process, as shown in Figure 1. Previous research that alleviates these errors lies in two main paradigms: 1) Answer aggregation across multiple rationales. They utilize majority voting (Wang et al., 2022) or deploy a verifier to score on each rationale (Li et al., 2023). However, these methods do not directly address errors in the reasoning process, undermining the reliability of their outcomes. 2) Intermediate step correction. This line of works decomposes the reasoning path into reasoning steps and applies self-correction on each step (Weng et al., 2022; Ling et al., 2024; Paul et al., 2023; Xie et al., 2023). Yet, recent research finds that, without external feedback, LLMs tend to modify reasoning steps regardless of their correctness (Huang et al., 2023; Hong et al., 2023).
@@
Previous works fail to address reasoning errors in intermediate steps, compromising the ability to conduct systematic reasoning. To mitigate this issue, we embrace the principle of deductive reasoning (Clark, 1969; Johnson-Laird, 1999; 2010). In deductive reasoning, every step logically follows its premises, where a deducible reasoning step is termed a logical consequence (Dinkmeyer, 1976; Hanson, 1997). A key attribute of logical consequence is that if the premises hold, the deducible reasoning step is true, suggesting a correct outcome. Inspired by this attribute, we propose to navigate CoT towards a more deducible path.
@@
Nonetheless, challenges arise when introducing the principle of deductive reasoning into CoT reasoning without changing the standard prompt paradigm and the parameters of LLMs. 1) Navigation on CoT reasoning. Since LLMs cannot always conduct correct deductive reasoning, they have to explore the potential reasoning space and choose those reasoning steps that are more likely to be deducible, which brings the trade-off between exploration and exploitation (Donoso et al., 2014; Dasgupta et al., 2019). 2) Verification of deducibility. On one hand, previous research shows it is hard for LLMs to detect reasoning errors (Huang et al., 2023; Hong et al., 2023). On the other hand, symbolic reasoning engines (Cavada et al., 2014; Li et al., 2018) can reliably verify the correctness. However, transferring natural language to symbolic language without losing generality remains an unsolved problem in machine reasoning.
@@
Confronted with these challenges, we propose Deductive Beam Search (DBS), adaptable to all models and settings. The overview of DBS is shown in Figure 2. For the trade-off challenge, we decompose the reasoning process into reasoning steps and incorporate step-wise beam search. In terms of the verification challenge, we propose a deductive verifier, which takes a reasoning step and its premises as inputs and outputs a deductive score, evaluating the logical coherence between them. Specifically, LLM samples a list of potential reasoning steps to explore. Then, our deductive verifier exploits by selecting steps that are more deducible. To train an effective verifier, we propose a scalable way of synthesizing fine-grained and diverse deductive reasoning errors without human annotation. Initially, the verifier is trained to verify heuristically synthesized wrong steps with typical reasoning error patterns. Subsequently, we ask LLMs to generate reasoning steps where false ones detected by our verifier serve as hard negatives. These hard negatives are adopted to train a deductive verifier with model feedback.
@@ 2.1 Multi-Step Chain-of-Thought Reasoning
$\Pr{}_{LM}(\textbf{r}_{[t]}|\textbf{q})=\Pr{}_{LM}(\textbf{r}_{1}|\textbf{q})\times\prod_{i=1}^{t-1}{\Pr{}_{LM}(\textbf{r}_{i+1}|\textbf{q},\textbf{r}_{[i]})}=\prod_{i=0}^{t-1}{\Pr{}_{LM}(\textbf{r}_{i+1}|\textbf{q},\textbf{r}_{[i]})}.$ (2)
@@ 2.2 Step-wise Beam Search
$Y_{[i]}=\operatorname*{arg\,max}_{\textbf{y}_{[i]}^{1},\textbf{y}_{[i]}^{2},...,\textbf{y}_{[i]}^{m}}\sum_{k\in[m]}\Phi(\textbf{y}_{[i]}^{k}).$ (3)
@@
$R_{[i]}=\operatorname*{arg\,max}_{\textbf{r}_{[i]}^{1},\textbf{r}_{[i]}^{2},...,\textbf{r}_{[i]}^{m}}\sum_{k\in[m]}\Psi(\textbf{r}_{[i]}^{k}).$ (4)
Combining multi-step CoT reasoning with step-wise beam search balances exploration and exploitation in reasoning tasks. However, confidence scores from language models cannot verify logical consequence between a reasoning step and its premises. To tackle this problem, we propose to constrain the step-wise beam search with deductive scores.
@@ 2.3 Deductive Verification Constrained Beam Search
$s_{j}=\textbf{s}([\textbf{r}_{[i-1]},\textbf{r}_{j}])=\textbf{s}(\textbf{r}_{[i-1]})\times\Pr{}_{f}(\textbf{r}_{j}|\textbf{q},\textbf{r}_{[i-1]})=\prod_{k=1}^{i}{\Pr{}_{f}(\textbf{r}_{k}|\textbf{q},\textbf{r}_{[k-1]}}),$ (5)
@@ 3 Deductive Verifier
As stated above, a deductive verifier evaluates whether the reasoning step can be deduced from previous contexts, which resembles a natural language inference (NLI) task. Thus, we use deberta-v3-large (He et al., 2021), which achieves the best performance across various NLI benchmarks despite its small size, as the backbone. A small scalar head is adopted to predict deductive scores based on embedding the [CLS] token.
However, the difficulties of training a deductive verifier primarily reside in the training data quality and the training method. Changing one single token could lead to various errors, which is hard for any model to detect. Furthermore, the lack of high-quality false deductive reasoning step hinges the training of the verifier. To fully understand how LLMs make mistakes, we dive into the incorrect samples generated by LLMs. From the perspective of deductive reasoning, there are two main classes of reasoning errors: grounding errors and logic errors (Ling et al., 2024). Most grounding errors happening in the reasoning process can be detected by finding the contradiction between the context and the rationales, while the latter ones are illogical reasoning steps deduced from the previous context.
Context Type Not. Reasoning Step Error Reason
Randy has some money. He spent \$10 buying his lunch. He spent a quarter of the money he had left on an ice cream cone. If the ice cream cone cost \$5, what is the amount of money, in dollars, Randy had at first? Gold r Randy has 5*4=20 dollars left after buying lunch. -
Grounding $\textbf{r}_{1}^{\prime}$ Randy has 10*4=40 dollars left after buying lunch. Minor token-level error, hard for models to detect.
Logic $\textbf{r}_{2}^{\prime}$ At first, Randy had a sum of 20+10=30 dollars. Logic-level error caused by reversed steps, not following deductive reasoning.
Irrelevant $\textbf{r}_{3}^{\prime}$ He eats 65 black cookies from the cookie jar, with 1/2 * 130 = 65. Major error, completely incoherent with the context.
Table 1: Examples of heuristically synthesized false reasoning steps.
@@ 3.1 A General Deductive Verifier
where $m_{i}$ is the hyper-parameter controlling the margin and $\alpha_{i}$ weighs each loss.
@@ 3.2 Deductive Verifier with Model Feedback
In the first stage, we train a general deductive verifier, but the wrong samples synthesized heuristically are less diverse than the ones encountered during inference. To bridge the gap between synthesized data and real-world data, we use the verifier from stage 1 to detect false reasoning steps generated by an actual language model, where we choose Llama2-7b for the generation. The reason why we choose a relatively small language model for the generation is to maximize the diversity and the likelihood of generating incorrect steps.
To be concrete, given the verifier $f_{1}$ trained by stage 1, we feed context c into the LLM and sample 10 reasoning steps. Then, these steps are scored and ranked by $f_{1}$. From this ranking, we select the reasoning step that exhibits the most significant decrease in the deductive score, designating it as the hard negative sample. We replace $\textbf{r}_{1}^{\prime}$ with the generated hard negative sample as $\textbf{r}_{1}^{\prime\prime}$, keeping the original way of generating $\textbf{r}_{2}^{\prime}$ and $\textbf{r}_{3}^{\prime}$. Consequently, we continue training the verifier $f_{1}$ by Equation 6 with a smaller learning rate.
4 Experimental Setup
4.1 Reasoning Tasks
Symbolic Reasoning. We use the Coin Flip dataset (Wei et al., 2022b). The task is to determine which face of the coin is up after a series of operations.
4.2 Details
Language Models. We evaluate our method on models of various scales, including Llama2-7b, Llama2-13b, Llama2-70b (Touvron et al., 2023), and ChatGPT (gpt-3.5-turbo-instruct) (OpenAI, 2022). These models represent different levels of reasoning abilities. For the verifier, we choose deberta-v3-large as the backbone of our verifier. The training details are in Appendix B.
Baselines. To prove the effectiveness of DBS, we compare with greedy decoding algorithm (Jurafsky & Martin, ) and self-consistency (Wang et al., 2022). For other SOTA baselines, we choose SelfEval (Xie et al., 2023) and Deductive Verification (Ling et al., 2024), which do not update the parameters of LLMs. The former represents SOTA decoding algorithm, while the latter stands for methods utilizing novel procedure design to conduct deductive reasoning. Since a full-scale experiment requires excessive token cost due to the extensive search and verification of these methods, we provide results on the GSM8K dataset.
@@ 4.2 Details
Table 2: The result comparison on arithmetic reasoning, commonsense reasoning, and symbolic reasoning tasks. The results represent accuracy (%) on each dataset. Bold indicates best results and underline indicates second bests.
@@ 5 Main Result
Table 2 demonstrates the overall performance of the methods. We compare DBS with baselines under two paradigms: single chain setting and multiple chain setting. In multiple chain setting, the generated outcomes are integrated with self-consistency. Table 3 presents a comparative analysis of our approach against SOTA baselines.
5.1 Effectiveness
As shown in Table 2, DBS improves the performance across models of different scales and diverse reasoning tasks. For the single chain setting, the improvement is substantial. On arithmetic reasoning tasks, taking GSM8K as an example, we observe an increase from 7.6% to 16.6% across models of various scales. Specifically, with Llama2-7b and gpt-3.5-turbo-instruct, DBS yields improvements of 9.2% and 7.0%, respectively, affirming the effectiveness of our proposed strategy. On commonsense reasoning tasks and symbolic reasoning tasks, we can expect an average increase of 2.5%/2.0% on models of all scales.
5.2 Comparison with Current Solutions
@@
Llama2-7b SelfEval 21.8 50M 1
SelfEval + SC 24.2 10
DV 10.5 372M 1
DV + SC 13.2 10
DBS 31.2 4M 1
DBS + SC 32.1 10
ChatGPT SelfEval 71.3 12M 1
SelfEval + SC 74.7 10
DV 68.2 409M 1
DV + SC 83.3 10
DBS 75.9 5M 1
DBS + SC 83.5 10
Table 3: Comparison of between DBS and different reasoning methods.
Table 3 compares our decoding strategy with previous SOTA reasoning strategies. The comparative results, grounded in accuracy and token cost metrics, substantiate our approach’s effectiveness and token efficiency. Notably, the self-evaluate pattern performs worse when the scale of the LLM drops. Moreover, it consumes an excessive amount of tokens during evaluation. In contrast, DBS enhances the performance by approximately 10% on Llama2-7b and performs better than the baselines on gpt-3.5-turbo-instruct across the paradigms of single and multiple reasoning chains. Furthermore, it consumes much fewer input and output tokens. Consuming 80 times less tokens, DBS outperforms DV on gpt-3.5-turbo-instruct.